p = 0.275 p^2 = 0.275 × 0.275 = 0.07562 q = 0.725 q^2 = 0.725 × 0.725 = 0.5256 2pq = 2 × 0.275 × 0.725 = 0.3987 p^2 + 2pq + q^2 = 0.99992 = approx. 1

What are the phenotypic ratio’s based on this Hardy Weinberg equilibrium? Please help me out thank you
Answer #1

Answer:

Dominant phenotype = p^2 + 2pq = 0.07562 + 0.3987 = 0.47432

Recessive phenotype = q^2 = 0.5256

Phenotypic rati = 1:1